OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x^3 + 4 and g(x) = x-4 cubed

OpenStudy (anonymous):

\[f(x)=x^3+4\] \[\sqrt{x-4}\]

OpenStudy (anonymous):

\[g(x)=\sqrt{x-4}\]